OpenStudy (anonymous):

Solve for y: 3x + 5y = 15

OpenStudy (anonymous):

You want to isolate the y term 3x + 5y = 15 First subtract the 3x - 3x -3x ------------- 5y = -3x + 15 then divide by 5 -------------- 5 y = -3/5x +3